log5x + log25x = 5
Answer by venugopalramana(3286)

log5x + log25x = 5
HOPE YOU MEAN ALL LOGS ARE TO BASE 10.
log(5x*25x)=5
log(125x^2)=5
125x^2 = 10^5
X^2=10^5/125=800
X = 20 SQRT(2) = 2828